# 2D or not 2D? Exploring 3D relativistic magnetic reconnection dynamics with highly accurate numerical simulations

### v1

@@ 3.2 Pressure-balanced configuration
At $t=15\tau_{c}$ (Figs. 7 and 7), the thermal-pressure is higher on the current sheet than in the upstream plasma due to the initial equilibrium configuration. Some signs of reconnection along the current sheet are visible (see also the $z=0$ cut in the mid-right panel of Fig. 6 at the corresponding time). The plasma $\beta$ on the current sheet is very high ($\beta_{\rm min}\simeq\beta_{\rm max}\simeq 10^{2}$ at $t=0$, and $\beta_{\rm min}\simeq 10^{-1}$ while $\beta_{\rm max}\simeq 10^{7}$ at $t=15\tau_{c}$). High $\beta$ values on the current sheet favour the onset of pressure-driven modes: as $\beta$ increases, the destabilizing effect of local pressure gradients becomes stronger relative to the magnetic tension that stabilizes the system, making it easier for perturbations to grow in regions of unfavourable curvature (Biskamp_1986). In our configuration, the magnetic tension is weak due to the absence of a guide field component ($B_{z}=0$) in the initial setup. Consequently, pressure-driven instabilities can compete with or alter other dynamics, such as reconnection, by rapidly acting on the formed plasmoids, which are destroyed before they can coalesce. This dynamics is suggested also by the Fourier spectrum at this stage, which is dramatically different from that of model ff3d at similar times (compare Fig. 7 with Fig. 3). Fig. 7 shows the excitation of the $m=1$ mode together with a broad range of low- to intermediate-$z$ wavelengths, as evidenced by the dominance of modes $\mathcal{M}_{1,n}$ with $n\lesssim 30$, which are a signature of emerging pressure-driven modes (Strauss_1981; Biskamp_MagRec). The suppression of the second linear phase, associated with the linear growth of the plasmoid coalescence instability in the upper panel of Fig. 5, along with the irregular morphology of plasmoids visible in the lower-right panel of Fig. 6, corroborates this dynamics. The absence of a guide field in the $z$-direction in the 3D setup is crucial for establishing this behaviour, along with the role of dimensionality. In fact, in the 2D pressure-balanced configuration, this effect cannot occur, since the lack of $z$-dependence precludes the growth of modes with $n>0$, thus leading the planar symmetry to mimic the stabilizing effect of a guide field.
@@ 3.3 Non-ideal regions
To compare the relative efficiency of magnetic reconnection in our models we quantify their deviation from the ideal RMHD regime by focusing on the evolution of the Lorentz invariant $E^{2}-B^{2}$. This quantity is strictly negative in the case of ideal electric fields $\mathbf{E}=-\mathbf{v}\times\mathbf{B}$, but it typically becomes positive in between merging plasmoids, where strong resistive electric fields are produced (Sironi_2022).
@@ 4 Summary and conclusions
In this work we investigated relativistic magnetic reconnection triggered by the ideal tearing mode instability within the framework of resistive special-relativistic MHD. Our study extends previous efforts that focused primarily on 2D configurations (e.g., Landi_etal2015; Tenerani_2015; DelZanna_2016) by performing, for the first time, fully 3D ResRMHD simulations of ideal-tearing–driven current-sheet disruption in both force-free and pressure-balanced equilibria. Each setup was run in 3D domains and in corresponding 2D reference configurations to isolate the effects of dimensionality. The simulations were carried out with the GPU-accelerated version of the PLUTO code (Mignone_2007; Rossazza_GPU_2025), using the high-order numerical scheme of Mignone_Berta_2024.
@@
In the force-free configuration (i.e., with a non-zero $B_{z}$ component), the tearing instability develops similarly in both 2D and 3D. In 2D, the system follows the well studied dynamics (e.g., Miranda_etal2018; Mignone_etal_2019; Mignone_Berta_2024): hierarchical plasmoid-chain reconnection drives energy release but saturates as magnetic energy remains stored in plasmoids, quenching conversion efficiency (Sironi_Spitkovsky_2014; Sironi_Petropoulou_Giannios_2015). In 3D, we observe the onset of ideal tearing in agreement with linear theory (Pucci_Velli_2014; DelZanna_2016), followed by parasitic plasmoid coalescence in close resemblance with the 2D reference case. At later stages, the additional degree of freedom (the $z$-direction) enables a richer dynamics characterized by secondary instabilities (i.e., flux-rope kinking, and thickening of the layer) which excite power at progressively smaller scales along the guide-field direction. This process breaks the planar coherence of the layer, triggers a turbulent cascade, which fosters field dissipation. This interplay produces a longer-lasting reconnection and ultimately leads to greater magnetic-energy dissipation than in 2D.
@@
The 2D pressure-balanced reconnection dynamics is qualitatively similar to that of the corresponding 2D force-free model (Landi_etal2015), though the ideal tearing instability grows at a slower rate. On the other hand, in the 3D pressure-balanced model with an initial null $B_{z}$ component, plasmoid-mediated reconnection is strongly suppressed compared to the 2D case and to force-free models. While the initial tearing instability develops similarly in both 2D and 3D pressure-balanced cases, the 3D model lacks the characteristic second exponential growth associated with the onset of the parasitic plasmoid coalescence instability (Loureiro_2007; Loureiro_Uzdensky_2016) triggered by tearing, along with a turbulent morphology of forming plasmoids. The forming plasmoids are disrupted by rapidly growing pressure-driven modes before they can merge, preventing the characteristic rise-and-saturation of the reconnected magnetic field profile, and the formation of a main magnetic island. This leads to a reduced reconnection efficiency, evidenced by weaker current spikes, and limited conversion of magnetic energy. A similar dynamics is observed in recent 3D first-principles closed-box simulations (e.g., Werner_Uzdensky_2021). Local plasma conditions on the current sheet promote the growth of pressure-driven modes, which, in the absence of a stabilizing $B_{z}$ component providing magnetic tension, grow rapidly along the $z$-direction and disrupt forming plasmoids. In 2D, such suppression cannot occur, as the lack of $z$-dependence prevents the growth of modes along $z$ and mimic the stabilizing effect of a guide field.
@@
Comparison of the 3D equilibria to their respective 2D reference cases shows that the presence of a $z$-dependence in the configuration, the 3D magnetic field topology, and the local plasma conditions critically determine the dynamics and efficiency of reconnection. The presence of a $B_{z}$ component in 3D enables ideal tearing to form plasmoids that persist long enough to merge, thereby triggering a bursty reconnection event. In 2D, the lack of a $z$-dependence and a $z$-direction forces the system into planar symmetry, effectively mimicking the presence of a guide field and preventing the emergence of parasitic modes and their associated dynamics. While 2D reconnection can be bursty due to stochastic creation, movement, and merging of plasmoids, these events tend to invariably lead toward a unique final state. In 3D, current-sheet evolution is not simply classical reconnection with additional perturbations. When considering a non-zero $B_{z}$ component, at least three main instabilities (and in reality, even more) contribute to the dynamics: the primary one is tearing, followed by plasmoid coalescence and the flux-rope kink instability, all fuelled by magnetic energy (Zhang_Kink_2021; Zhang_2023_kink_Flux_rope; Sironi_2025_RecReview). In contrast, when $B_{z}$ is zero, other instabilities not necessarily powered by magnetic energy -such as pressure-driven modes- may arise and concur in affecting the development of the primary ideal-tearing event. Thus, in 3D, the reconnection problem effectively requires the understanding of the dynamical behaviour of thin current sheets governed by the non-linear interaction of several instabilities, which may act simultaneously and either compete with or reinforce one another. The outcome of this interplay depends sensitively on the ambient plasma, subtle configuration details, and even stochastic effects (Werner_Uzdensky_2021), and it is crucial for determining the overall evolution and efficiency of reconnection-something that can only be assessed within a fully 3D framework.
@@
Several directions for future work emerge from our study. First, it is crucial to assess the sensitivity of reconnection dynamics to the initial and boundary conditions, in particular by exploring open boundary conditions and the effect of a guide field in pressure-balanced equilibria. Secondly, recent results from particle-in-cell (PIC) simulations (Selvi_etal_2023; Moran_2025) suggest that the fundamental properties governing magnetic-field dissipation in current sheets can be captured by an effective resistivity formulation. This approach locally enhances magnetic energy dissipation and favours the onset of fast reconnection (Bugli_etal_2024). Applying collisionless resistivity prescriptions to 3D ResRMHD reconnection models is therefore essential for resolving collisionless magnetic reconnection physics while achieving scale separations that are prohibitive for fully kinetic models. Such studies are fundamental for developing reconnection models at larger scales, with potential applications including disentangling the coherent radio emission from the Crab pulsar, other young energetic pulsars, and millisecond pulsars produced in the magnetospheric current sheet beyond the light cylinder (Philippov_Uzdensky_Spitkovsky_Cerutti_2019), as well as the dissipation of the striped pulsar wind in pulsar magnetospheres (Cerutti_2017_Striped; Cerutti_2020_Striped). Moreover, coupling the findings of our simulations with large-scale models represents a crucial step toward properly isolating observable features triggered by magnetic reconnection events in simulations of black hole or neutron star magnetospheres (Tchekhovskoy_2013), AGN jets (Mattia2023; Mattia_etal_2024), coronae (Liska_2022), accretion disks (Ripperda_2022_BH_Flares), as well as collapsing or merging objects (Most_2024).
@@ Acknowledgements
The data underlying this article will be shared on reasonable request to the corresponding author. The figures presented in this paper have been generated using the PyPLUTO package (Mattia_etal_2025) and VisIt (VisIt).

### v2

@@ 3.2 Pressure-balanced configuration
At $t=15\tau_{c}$ (Figs. 7 and 7), the thermal-pressure is higher on the current sheet than in the upstream plasma due to the initial equilibrium configuration. Some signs of reconnection along the current sheet are visible (see also the $z=0$ cut in the mid-right panel of Fig. 6 at the corresponding time). The plasma $\beta$ on the current sheet is very high ($\beta_{\rm min}\simeq\beta_{\rm max}\simeq 10^{2}$ at $t=0$, and $\beta_{\rm min}\simeq 10^{-1}$ while $\beta_{\rm max}\simeq 10^{7}$ at $t=15\tau_{c}$). High $\beta$ values on the current sheet favour the onset of pressure-driven modes: as $\beta$ increases, the destabilizing effect of local pressure gradients becomes stronger relative to the magnetic tension that stabilizes the system, making it easier for perturbations to grow in regions of unfavourable curvature (Biskamp, 1986). In our configuration, the magnetic tension is weak due to the absence of a guide field component ($B_{z}=0$) in the initial setup. Consequently, pressure-driven instabilities can compete with or alter other dynamics, such as reconnection, by rapidly acting on the formed plasmoids, which are destroyed before they can coalesce. This dynamics is suggested also by the Fourier spectrum at this stage, which is dramatically different from that of model ff3d at similar times (compare Fig. 7 with Fig. 3). Fig. 7 shows the excitation of the $m=1$ mode together with a broad range of low- to intermediate-$z$ wavelengths, as evidenced by the dominance of modes $\mathcal{M}_{1,n}$ with $n\lesssim 30$, which are a signature of emerging pressure-driven modes (Strauss, 1981; Biskamp, 2000). The suppression of the second linear phase, associated with the linear growth of the plasmoid coalescence instability in the upper panel of Fig. 5, along with the irregular morphology of plasmoids visible in the lower-right panel of Fig. 6, corroborates this dynamics. The absence of a guide field in the $z$-direction in the 3D setup is crucial for establishing this behaviour, along with the role of dimensionality. In fact, in the 2D pressure-balanced configuration, this effect cannot occur, since the lack of $z$-dependence precludes the growth of modes with $n>0$, thus leading the planar symmetry to mimic the stabilizing effect of a guide field.
@@ 3.3 Non-ideal regions
To compare the relative efficiency of magnetic reconnection in our models we quantify their deviation from the ideal RMHD regime by focusing on the evolution of the Lorentz invariant $E^{2}-B^{2}$. This quantity is strictly negative in the case of ideal electric fields $\mathbf{E}=-\mathbf{v}\times\mathbf{B}$, but it typically becomes positive in between merging plasmoids, where strong resistive electric fields are produced (Sironi, 2022).
@@ 4 Summary and conclusions
In this work we investigated relativistic magnetic reconnection triggered by the ideal tearing mode instability within the framework of resistive special-relativistic MHD. Our study extends previous efforts that focused primarily on 2D configurations (e.g., Komissarov et al., 2007; Landi et al., 2015; Tenerani et al., 2015b; Del Zanna et al., 2016) by performing, for the first time, fully 3D ResRMHD simulations of ideal-tearing–driven current-sheet disruption in both force-free and pressure-balanced equilibria. Each setup was run in 3D domains and in corresponding 2D reference configurations to isolate the effects of dimensionality. The simulations were carried out with the GPU-accelerated version of the PLUTO code (Mignone, 2007; Rossazza et al., 2025), using the high-order numerical scheme of Mignone et al. (2024).
@@
In the force-free configuration (i.e., with a non-zero $B_{z}$ component), the tearing instability develops similarly in both 2D and 3D. In 2D, the system follows the well studied dynamics (e.g., Miranda-Aranguren et al., 2018; Mignone et al., 2019, 2024): hierarchical plasmoid-chain reconnection drives energy release but saturates as magnetic energy remains stored in plasmoids, quenching conversion efficiency (Sironi and Spitkovsky, 2014; Sironi et al., 2015). In 3D, we observe the onset of ideal tearing in agreement with linear theory (Pucci and Velli, 2014; Del Zanna et al., 2016), followed by parasitic plasmoid coalescence in close resemblance with the 2D reference case. At later stages, the additional degree of freedom (the $z$-direction) enables a richer dynamics characterized by secondary instabilities (i.e., flux-rope kinking, and thickening of the layer) which excite power at progressively smaller scales along the guide-field direction. This process breaks the planar coherence of the layer, triggers a turbulent cascade, which fosters field dissipation. This interplay produces a longer-lasting reconnection and ultimately leads to greater magnetic-energy dissipation than in 2D.
@@
The 2D pressure-balanced reconnection dynamics is qualitatively similar to that of the corresponding 2D force-free model (Landi et al., 2015), though the ideal tearing instability grows at a slower rate. On the other hand, in the 3D pressure-balanced model with an initial null $B_{z}$ component, plasmoid-mediated reconnection is strongly suppressed compared to the 2D case and to force-free models. While the initial tearing instability develops similarly in both 2D and 3D pressure-balanced cases, the 3D model lacks the characteristic second exponential growth associated with the onset of the parasitic plasmoid coalescence instability (Loureiro et al., 2007; Loureiro and Uzdensky, 2016) triggered by tearing, along with a turbulent morphology of forming plasmoids. The forming plasmoids are disrupted by rapidly growing pressure-driven modes before they can merge, preventing the characteristic rise-and-saturation of the reconnected magnetic field profile, and the formation of a main magnetic island. This leads to a reduced reconnection efficiency, evidenced by weaker current spikes, and limited conversion of magnetic energy. A similar dynamics is observed in recent 3D first-principles closed-box simulations (e.g., Werner and Uzdensky, 2021). Local plasma conditions on the current sheet promote the growth of pressure-driven modes, which, in the absence of a stabilizing $B_{z}$ component providing magnetic tension, grow rapidly along the $z$-direction and disrupt forming plasmoids. In 2D, such suppression cannot occur, as the lack of $z$-dependence prevents the growth of modes along $z$ and mimic the stabilizing effect of a guide field.
@@
Comparison of the 3D equilibria to their respective 2D reference cases shows that the presence of a $z$-dependence in the configuration, the 3D magnetic field topology, and the local plasma conditions critically determine the dynamics and efficiency of reconnection. The presence of a $B_{z}$ component in 3D enables ideal tearing to form plasmoids that persist long enough to merge, thereby triggering a bursty reconnection event. In 2D, the lack of a $z$-dependence and a $z$-direction forces the system into planar symmetry, effectively mimicking the presence of a guide field and preventing the emergence of parasitic modes and their associated dynamics.
@@
While 2D reconnection can be bursty due to stochastic creation, movement, and merging of plasmoids, these events tend to invariably lead toward a unique final state. In 3D, current-sheet evolution is not simply classical reconnection with additional perturbations. When considering a non-zero $B_{z}$ component, at least three main instabilities (and beyond a single fluid description, even more, such as the Relativistic Drift Kink Instability) contribute to the dynamics: the primary one is tearing, followed by plasmoid coalescence and the flux-rope kink instability, all fueled by magnetic energy (Barkov and Komissarov, 2016; Zhang et al., 2021b, a; Sironi et al., 2025). In contrast, when $B_{z}$ is zero, other instabilities not necessarily powered by magnetic energy -such as pressure-driven modes- may arise and concur in affecting the development of the primary ideal-tearing event. Thus, in 3D, the reconnection problem effectively requires the understanding of the dynamical behaviour of thin current sheets governed by the non-linear interaction of several instabilities, which may act simultaneously and either compete with or reinforce one another. The outcome of this interplay depends sensitively on the ambient plasma, subtle configuration details, and even stochastic effects (Werner and Uzdensky, 2021), and it is crucial for determining the overall evolution and efficiency of reconnection-something that can only be assessed within a fully 3D framework.
@@
Several directions for future work emerge from our study. First, it is crucial to assess the sensitivity of reconnection dynamics to the initial and boundary conditions, in particular by exploring open boundary conditions and the effect of a guide field in pressure-balanced equilibria. Secondly, recent results from particle-in-cell (PIC) simulations (Selvi et al., 2023; Moran et al., 2025) suggest that the fundamental properties governing magnetic-field dissipation in current sheets can be captured by an effective resistivity formulation. This approach locally enhances magnetic energy dissipation and favours the onset of fast reconnection (Bugli et al., 2025; Ripperda et al., 2026).
Applying collisionless resistivity prescriptions to 3D ResRMHD reconnection models is therefore essential for resolving collisionless magnetic reconnection physics while achieving scale separations that are prohibitive for fully kinetic models. Such studies are fundamental for developing reconnection models at larger scales, with potential applications including disentangling the coherent radio emission from the Crab pulsar, other young energetic pulsars, and millisecond pulsars produced in the magnetospheric current sheet beyond the light cylinder (Baty et al., 2013; Philippov et al., 2019), as well as the dissipation of the striped pulsar wind in pulsar magnetospheres (Cerutti and Philippov, 2017; Cerutti et al., 2020). Moreover, coupling the findings of our simulations with large-scale models represents a crucial step toward properly isolating observable features triggered by magnetic reconnection events in simulations of black hole or neutron star magnetospheres (Tchekhovskoy et al., 2013), AGN jets (Mattia et al., 2023), coronae (Liska et al., 2022), accretion disks (Ripperda et al., 2022a), as well as collapsing or merging objects (Most et al., 2024).
@@ Acknowledgements
The data underlying this article will be shared on reasonable request to the corresponding author. The figures presented in this paper have been generated using the PyPLUTO package (Mattia et al., 2025) and VisIt (Childs et al., 2012).